In a box there are 7 balls, of which 3 are red and 4 are white. Three balls will be withdrawn at random. Find the probability that at least one ball will be red.
Thank you for your help. Best regards

Answer by stanbon(75887)

In a box there are 7 balls, of which 3 are red and 4 are white. Three balls will be withdrawn at random. Find the probability that at least one ball will be red.
----------------
P(at least one ball is red) = 1 - P(no ball is red)
--------------
P(at least one) = 1 - [4C3/7C3] = 1 - 4/35 = 31/35
=======================================================
Cheers,
Stan H.